OpenStudy (anonymous):

Find the value of x if the area of the triangle is 280 in squared?

OpenStudy (anonymous):



OpenStudy (jdoe0001):

area of a triangle = \(\bf \cfrac{1}{2} \times base \times height\) 

OpenStudy (jdoe0001):

so we know the area is 280, we know the base is 70 inches, thus \(\bf Area = \cfrac{1}{2} \times base \times height\\ 280 = \cfrac{1}{2} \times 76 \times height\) solve for "height"

OpenStudy (anonymous):

I got 7.37 for height, :/ but that's not one of the options

OpenStudy (jdoe0001):

ohh smokes, yes, I put 76, but it's 70

OpenStudy (jdoe0001):

\(\bf Area = \cfrac{1}{2} \times base \times height\\ 280 = \cfrac{1}{2} \times 70 \times height\)

OpenStudy (jdoe0001):

my bad typo

OpenStudy (anonymous):

haha no problem. I got 8, is that correct?

OpenStudy (jdoe0001):

yes